In the final stages of a moon landing, a lunar module descends under retro-thrust of its descent engine to a height h = 3.7 m above the lunar surface where it has a downward velocity of 4m/s. Calculate the impact velocity with the moon’s surface if the engine is cut off at this point and lunar gravity is 1/6 of the earth’s gravity.
1
Expert's answer
2019-08-14T09:43:56-0400

Formulae adopted is

"V_{f}^2=V_{i}^2+2\times g\times d"

Where g is acceleration due to gravity on moon which is 1/6 times of the earth, thus

g=9.81/6=1.6 m/s2

Vi=4m/s

d=3.7meter

Thus,

"V_{f}^2=4^2+2\times1.6\times3.7"

Thus Vf=5.27m/s